# GVDepth: Zero-Shot Monocular Depth Estimation for Ground Vehicles based on Probabilistic Cue Fusion

### v1

@@ Abstract
Generalizing metric monocular depth estimation presents a significant challenge due to its ill-posed nature, while the entanglement between camera parameters and depth amplifies issues further, hindering multi-dataset training and zero-shot accuracy. This challenge is particularly evident in autonomous vehicles and mobile robotics, where data is collected with fixed camera setups, limiting the geometric diversity. Yet, this context also presents an opportunity: the fixed relationship between the camera and the ground plane imposes additional perspective geometry constraints, enabling depth regression via vertical image positions of objects. However, this cue is highly susceptible to overfitting, thus we propose a novel canonical representation that maintains consistency across varied camera setups, effectively disentangling depth from specific parameters and enhancing generalization across datasets. We also propose a novel architecture that adaptively and probabilistically fuses depths estimated via object size and vertical image position cues. A comprehensive evaluation demonstrates the effectiveness of the proposed approach on five autonomous driving datasets, achieving accurate metric depth estimation for varying resolutions, aspect ratios and camera setups. Notably, we achieve comparable accuracy to existing zero-shot methods, despite training on a single dataset with a single-camera setup. Project website: https://gvdepth.github.io/
@@ 1 Introduction
In addition to the environmental domain gap, MDE is particularly sensitive to the domain gap induced by the perspective geometry parameters of the camera system. This issue is present in most supervised [11, 24, 1, 51, 34] and self-supervised methods [55, 52, 14, 15, 54, 53, 42] that are trained on a dataset collected with a single camera system. Models tend to overfit to this perspective geometry bias, leading to significant performance degradation when inferring depth with different cameras. Unfortunately, training a MDE model on data with a wide distribution of perspective geometries is far from trivial. Due to the inherent ambiguity of the MDE problem and entanglement of depth and camera parameters [35], many methods resort to relative depth estimation with unknown shift and scale [8, 43, 36, 48, 49, 21, 46]. Recently, zero-shot metric depth estimation has been achieved with various approaches, including embedding [16], estimating [3, 35] or disentangling the camera parameters [50]. While these methods present impressive results, they require training on data acquired from multiple camera systems, which is not always avaialable or cost-effective.
@@
One specific area with numerous applications across varying fields is MDE for ground-based vehicles and robots. In this context, an additional perspective geometry constraint arises due to the fixed position of the camera relative to the ground plane. This constraint induces an additional depth cue: the vertical image position of the ground-contact point, which has been shown to be a dominant cue in autonomous driving scenarios [10]. The vertical image position cue is closely related to the depth of the ground plane that has recently been leveraged to deliver State-of-the-Art (SotA) depth estimation accuracy in both supervised [47] and self-supervised [42, 4] settings. However, these methods demonstrate limited generalization results, as they are not designed with intention to enable zero-shot estimation.
@@
These findings motivate us to utilize the vertical image position cue and the inherent planarity of the scene in order to enable accurate depth estimation and improve generalization for arbitrary intrinsic and extrinsic camera parameters. Contrary to the object size cue, widely used by the aforementioned zero-shot methods [16, 3, 35, 50], we argue that the vertical image position presents a cue that is more fundamentally rooted in the perspective geometry configuration of the scene. However, without proper modeling this cue is highly susceptible to overfitting to the specific camera system present in the training data [23, 33].
@@
To leverage this cue, while enabling zero-shot generalization, in this paper we propose a method to ensure its consistency for arbitrary camera setups. Specifically, we introduce a novel intermediate depth representation termed Vertical Canonical Representation. Rather than directly regressing depth values, our model predicts vertical image positions of the orthogonal projections of points onto the ground plane, as visualized in Fig. 1. By utilizing the known depth of the ground plane, this approach allows the mapping between the canonical representation and depth space to adapt with varying camera parameters, effectively disentangling the vertical position cue from specific perspective geometry. To further address potential errors in certain regions, we introduce an architecture that adaptively fuses depth predictions from both vertical position and object size cues, guided by dynamically estimated uncertainties.
We conduct comprehensive evaluation of zero-shot MDE in autonomous driving scenarios, thoroughly analyzing the robustness and accuracy of depth estimation with different cues. We demonstrate that our approach is the first to achieve accurate out-of-distribution metric depth estimation when trained solely on a single dataset captured by a single vehicle-camera system. This marks a significant advancement through enabling generalizable metric MDE in data-scarce and specialized scenarios, such as mobile robotics.
@@ 2 Related Work
Domain-agnostic Monocular Depth Estimation. The inherent ambiguity and interdependence between camera parameters and depth present challenges for training a robust MDE model that generalizes across a wide range of environments. As datasets are usually collected with a single camera system, models tend to act discriminatively and associate perspective geometry parameters with the pertaining environment. Relative depth estimation methods [8, 43, 36, 48, 49, 46, 21] bypass camera parameter dependency by estimating depth with an unknown scale. While these approaches achieve visually impressive results, their inability to estimate metric depth limits their practicality in real-world applications, often necessitating integration with additional sensors.
@@
Planar Constraints for Depth Estimation. Real-world scenes are often highly regular, featuring numerous planar structures that can serve as additional cues to guide depth estimation [38, 32]. A highly relevant application of MDE is in autonomous driving and mobile robotics, where a fixed relation between the camera system and the ground plane induces additional perspective geometry constraints. It has been shown that models heavily utilize the corresponding ground-contact cues, completely disregarding the object size cue [10], unless models are trained with random crops or more advanced data grafting [33]. Moreover, incorporating ground plane parameterization within the model enhances model accuracy and robustness, both in supervised [47, 23], and self-supervised training [4, 42].
@@ 3 Proposed Method
Preliminaries. In this work, we examine an extremely common application domain of MDE: ground vehicles or a robot with a fixed camera configuration. Assuming an ego coordinate system defined at the ground level, we consider a camera system parameterized by intrinsic matrix $\mathbf{K}$ and extrinsic matrix $\mathbf{R|t}$, which transforms points from camera to ego coordinate frame:
$\mathbf{K}=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix},\mathbf{[R|t]}=\begin{bmatrix}1&0&0&0\\ 0&\cos\theta&-\sin\theta&-h\\ 0&\sin\theta&\cos\theta&0\\ \end{bmatrix},$ (1)
with $\theta,h$ representing camera pitch and height, respectively. We disregard the camera yaw and roll in our parameterization; yaw has no effect on our canonical transformations, while roll is typically negligible in real-world setups.
Problem Statement. We aim to learn a dense mapping $f:(\mathbf{I},\bm{\mathcal{C}})\mapsto\mathbf{D}$, with $\mathbf{I}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}\in\mathbb{R}^{H\times W}$ being RGB image and a correspondending depth map, and $\mathcal{C}$ representing camera parameterization in Eq. 1. Most importantly, the learned mapping should be able to generalize to images acquired with arbitrary camera system. In contrast to existing zero-shot methods [35, 16, 3, 50], which rely on multiple datasets with varying camera parameters, we aim to learn this mapping using data acquired with a single camera system $\bm{\mathcal{C}}_{\text{Train}}$.
@@ 3.1 Depth Cues
For ground vehicles, a fixed position of the camera system relative to the ground introduces an additional perspective geometry constraint – known depth of the ground plane, which is constant for all images acquired by the same camera setup. For each pixel $(u,v)$, depth $d$ is calculated from the intersection of the camera rays $d\,\mathbf{K}^{-1}[u,v,1]^{T}$ and the ground plane $\mathbf{R}^{T}\mathbf{n}-h=0$, with $\mathbf{n}=(0,-1,0)$ being the upward ground normal. This constraint enables depth regression via object’s ground contact point:
$d=\frac{f_{y}h}{\left(H-c_{y}-y\right)\cos(\theta)-f_{y}\sin(\theta)},$ (3)
where $y=H-v$ represents vertical image position of the ground projection, as visualized in Fig. 2. Compared to the object size cue, which is entirely semantic, this cue implicitly encodes the known perspective geometry constraints of the scene, thereby inducing a valuable prior to the ill-posed MDE problem. However, it is highly reliant on accurate and consistent camera calibration.
@@ 3.2 Canonical Transforms
Metric3D [50] proposed a canonical transform disentangling the predicted depth and camera parameters, enabling the use of object size cue for arbitrary camera systems. To achieve this, depth is predicted in canonical space $\mathbf{C_{F}}$ with a canonical focal length $f_{c}$, and post-processed to metric depth $\mathbf{D_{F}}$ by applying Focal Canonical Transform $\text{FCT}_{\bm{\mathcal{C}}}(\cdot)$, where $\mathbf{D_{F}}=\text{FCT}_{\bm{\mathcal{C}}}(\mathbf{C_{F}})=\frac{f_{y}}{f_{c}}\mathbf{C_{F}}$, and $\bm{\mathcal{C}}=\{f_{y}\}$.
Besides the object size cue, in this paper we propose to leverage the vertical image position cue, effectively encoding the configuration of the scene’s perspective geometry. Unfortunately, this cue also induces entanglement between camera parameters and depth. To enhance cross-dataset generalization, we introduce an additional canonical transform, which enables consistent exploitation of the ground plane prior. Under this formulation, the model estimates a dense map $\mathbf{C_{Y}}$ representing vertical image position of the ground projection point. Metric depth $\mathbf{D_{Y}}$ is then obtained via Vertical Canonical Transform $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$, where $\text{VCT}_{\bm{\mathcal{C}}}(\mathbf{C_{Y}})$ corresponds to mapping $y\mapsto d$ obtainable from Eq. 3. Therefore, this transformation requires knowledge of camera parameters $\bm{\mathcal{C}}=\{f_{y},c_{y},h,\theta\}$. Crucially, the proposed canonical space remains consistent under variations of camera parameters, facilitating generalization to diverse scenarios.
The proposed VCT is related to the depth of the ground plane, frequently used to enhance MDE accuracy in autonomous driving scenarios [47, 4, 42]. However, these methods exploit the ground-plane constraint exclusively for road pixels, resulting in limited generalization capability.
@@ 3.4 Geometric Augmentation
Contrary to zero-shot methods trained on multiple datasets [35, 16, 3, 50], we aim to achieve similar generalization capabilities by training on a single camera system. However, without geometric diversity induced by varying camera parameters, model can simply disregard our proposed canonical transformations. Therefore, we apply geometric augmentations, namely random cropping and resizing, to simulate different camera perspectives. Crucially, compared to [35, 3, 50], we train our model with multiple resolutions, enabling generalization for arbitrary image sizes. A comparable augmentation strategy is suggested in [42, 18], aiming to enhance generalization for arbitrary cameras. However, as we will demonstrate, without the use of canonical transformations, this approach does not effectively transfer across different datasets.
@@ 3.5 Optimization
$\mathcal{L}_{\text{unc}}=$ $\,\mathbf{\Sigma_{F}}^{-1}\odot||\mathbf{D_{F}-D^{*}}||_{1}+\mathbf{\Sigma_{Y}}^{-1}\odot||\mathbf{D_{Y}-D^{*}}||_{1}$
@@ 4 Experiments
Datasets. We use a diverse set of autonomous driving datasets, each collected with a different camera setup. Specifically, we utilize an aggregation of KITTI [13], DDAD [15], DrivingStereo [45], Waymo [40] and Argoverse Stereo [6] datasets. We split the data into training and validation sets as in the official data splits, utilizing images only from frontal cameras. Due to the missing and sometimes inaccurate information, we calibrate extrinsic parameters, i.e., camera height and pitch, which are required for our canonical representation. Further details are provided in the supplementary materials.
@@
Evaluation details. We use established depth metrics: absolute relative error ($\mathrm{A.Rel}$), root mean squared error ($\mathrm{RMS}$), root mean squared log error ($\mathrm{RMS}_{\log})$, square relative error ($\mathrm{Sq.Rel}$) and accuracy $\mathrm{\delta}_{i}$ under a threshold $1.25^{i}$. Depth is evaluated in the metric range of [0 m, 80 m].
@@ 4.1 Evaluation of Zero-shot Accuracy
Comparison with State-of-the-Art. We perform extensive evaluation, comparing GVDepth with SotA domain-dependent and domain-agnostic methods. Results for KITTI and DDAD datasets are shown in Tab. 1. Both supervised [34, 51] and self-supervised [14, 54] domain-dependent methods perform well on datasets matching their training domain but overfit to environmental and geometric biases, leading to reduced accuracy in KITTI $\rightarrow$ DDAD transfer. Relative depth estimation models exhibit greater adaptability across domains; however, they require ground-truth scale and shift alignment, limiting their practical applicability. Recent zero-shot metric methods, such as ZeroDepth, UniDepth, and Metric3D, demonstrate strong generalization capabilities but rely on extensive, diverse training datasets, which are not always feasible to obtain. Furthermore, UniDepth and Metric3D suffer accuracy degradation when inferring on resolutions lower than their high-resolution training data. While resizing and padding can address this, these adjustments increase computational demands, limiting their suitability for resource-constrained systems. GVDepth delivers comparable results while being adaptable to arbitrary camera setups, aspect ratios, and image resolutions. Depending on the train-test dataset combination, our method ranks first or second across most metrics, despite training on only a single dataset and using a fraction of the data required by other zero-shot MDE methods. This is particularly impressive, as each dataset features unique environments and camera setups, demonstrating that our model relies on domain-agnostic perspective geometry cues rather than domain-specific semantic cues.
@@
Table 2: Comparison with methods that leverage known ground plane. Results for KITTI $\rightarrow$ DDAD zero-shot transfer. (†): Results taken from corresponding papers.
PlaneDepth [42] 32.7 14.0 0.529 37.9
GEDepth [47] 26.1 16.1 - -
Groco  [4] 42.4 15.4 - -
GVDepth 11.8 8.35 0.251 82.5
In Tab. 2 we compare GVDepth zero-shot accuracy with models leveraging the known ground plane for increased performance. GVDepth outperforms competing methods by a significant margin, even though it relies on very similar planar constraint. This highlights the superiority of our Vertical Canonical Representation, which leverages the ground-plane constraint across all image pixels, in contrast to methods like [42, 47, 4] that only improve performance for road pixels.

### v2

@@ Abstract
Generalizing metric monocular depth estimation presents a significant challenge due to its ill-posed nature, while the entanglement between camera parameters and depth amplifies issues further, hindering multi-dataset training and zero-shot accuracy. This challenge is particularly evident in autonomous vehicles and mobile robotics, where data is collected with fixed camera setups, limiting the geometric diversity. Yet, this context also presents an opportunity: the fixed relationship between the camera and the ground plane imposes additional perspective geometry constraints, enabling depth regression via vertical image positions of objects. However, this cue is highly susceptible to overfitting, thus we propose a novel canonical representation that maintains consistency across varied camera setups, effectively disentangling depth from specific parameters and enhancing generalization across datasets. We also propose a novel architecture that adaptively and probabilistically fuses depths estimated via object size and vertical image position cues. A comprehensive evaluation demonstrates the effectiveness of the proposed approach on five autonomous driving datasets, achieving accurate metric depth estimation for varying resolutions, aspect ratios and camera setups. Notably, we achieve comparable accuracy to existing zero-shot methods, despite training on a single dataset with a single-camera setup. Project website: https://unizgfer-lamor.github.io/gvdepth/
@@ 1 Introduction
In addition to the environmental domain gap, MDE is particularly sensitive to the domain gap induced by the perspective geometry parameters of the camera system. This issue is present in most supervised [11, 23, 1, 50, 33] and self-supervised methods [54, 51, 14, 15, 53, 52, 41] that are trained on a dataset collected with a single camera system. Models tend to overfit to this perspective geometry bias, leading to significant performance degradation when inferring depth with different cameras. Unfortunately, training a MDE model on data with a wide distribution of perspective geometries is far from trivial. Due to the inherent ambiguity of the MDE problem and entanglement of depth and camera parameters [34], many methods resort to relative depth estimation with unknown shift and scale [8, 42, 35, 47, 48, 19, 45]. Recently, zero-shot metric depth estimation has been achieved with various approaches, including embedding [16], estimating [3, 34] or disentangling the camera parameters [49]. While these methods present impressive results, they require large-scale training on data acquired from multiple camera systems, which is not always practical or cost-effective.
@@
One specific area with numerous applications across varying fields is MDE for ground-based vehicles and robots. In this context, an additional perspective geometry constraint arises due to the fixed position of the camera relative to the ground plane. This constraint induces an additional depth cue: the vertical image position of the ground-contact point, which has been shown to be a dominant cue in autonomous driving scenarios [10].
@@
These findings motivate us to utilize the vertical image position cue and the inherent planarity of the scene in order to enable accurate depth estimation and improve generalization for arbitrary intrinsic and extrinsic camera parameters. Contrary to the object size cue, widely used by the aforementioned zero-shot methods [16, 3, 34, 49], we argue that the vertical image position presents a cue that is more fundamentally rooted in the perspective geometry configuration of the scene. However, without proper modeling, this cue is highly susceptible to overfitting to the specific camera system present in the training data [22, 32, 21].
@@
To leverage this cue, while enabling training with diverse perspective geometries and zero-shot generalization, in this paper we propose a method to ensure its consistency for arbitrary camera setups. Specifically, we introduce a novel intermediate depth representation termed Vertical Canonical Representation. Rather than directly regressing depth values, our model predicts vertical image positions of the orthogonal projections of points onto the ground plane, as visualized in Fig. 1. By utilizing the known depth of the ground plane, this approach allows the mapping between the canonical representation and depth space to adapt with varying camera parameters, effectively disentangling the vertical position cue from specific perspective geometry. To further address potential errors in certain regions, we introduce an architecture that adaptively fuses depth predictions from both vertical position and object size cues, guided by dynamically estimated uncertainties.
We comprehensively evaluate our method’s zero-shot capability on five autonomous driving datasets, analyzing its robustness and accuracy across various cues. We demonstrate that our approach matches the accuracy of existing zero-shot methods, despite using a single datasets collected with a single camera system.
To summarize, we present the following contributions:
  • a novel canonical transformation that enables the exploitation of vertical image position cue across diverse perspective geometries, enhancing both training scalabilty and zero-shot generalization,
  • a model architecture for the probabilistic fusion of depth maps estimated using object size and vertical image position cues,
  • a thorough ablation of different depth cues, with insights previously unseen in the literature.
@@ 2 Related Work
Domain-agnostic Monocular Depth Estimation. The inherent ambiguity and interdependence between camera parameters and depth present challenges for training a robust MDE model that generalizes across a wide range of environments. As datasets are usually collected with a single camera system, models tend to act discriminatively and associate perspective geometry parameters with the pertaining environment. Relative depth estimation methods [8, 42, 35, 47, 48, 45, 19] bypass camera parameter dependency by estimating depth with an unknown scale. While these approaches achieve visually impressive results, their inability to estimate metric depth limits their applicability, often necessitating integration with additional sensors.
@@
Planar Constraints for Depth Estimation. Real-world scenes are often highly regular, featuring numerous planar structures that can serve as additional cues to guide depth estimation [37, 31]. A highly relevant application of MDE is in autonomous driving and mobile robotics, where a fixed relation between the camera system and the ground plane induces additional perspective geometry constraints. It has been shown that models heavily utilize the corresponding ground-contact cues, completely disregarding the object size cue [10]. Moreover, incorporating ground plane constraint within the model enhances overall model accuracy, both in supervised [46], and self-supervised training [4, 41]. However, these methods leverage the known perspective geometry constraints to enhance depth estimation only for road pixels, leading to limited generalization capabilities.
@@ 3 Proposed Method
[…]
$\mathbf{K}=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix},\quad^{\mathcal{C}}_{\mathcal{E}}\mathbf{R}=\begin{bmatrix}1&0&0\\ 0&\sin\theta&-\cos\theta\\ 0&\cos\theta&\sin\theta\\ \end{bmatrix}$, (1)
and $\mathbf{t}=[0,0,-h]^{T}$, with $(\theta,h)$ representing camera pitch and height, respectively. Transformation of points from ego to camera coordinate frame is then: $\mathbf{p}_{\mathcal{C}}=\prescript{\mathcal{C}}{\mathcal{E}}{\mathbf{R}}\mathbf{p}_{\mathcal{E}}+\mathbf{t}$. We disregard the camera yaw and roll in our parameterization; yaw has no effect on our canonical transformations, while roll is typically negligible in real-world setups.
Problem Statement. We aim to learn a dense mapping $f:(\mathbf{I},\bm{\mathcal{C}})\mapsto\mathbf{D}$, with $\mathbf{I}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}\in\mathbb{R}^{H\times W}$ being RGB image and a correspondending depth map, and $\bm{\mathcal{C}}$ representing camera parameterization in Eq. 1. Most importantly, the learned mapping should be able to generalize to images acquired with arbitrary camera system. In contrast to existing zero-shot methods [34, 16, 3, 49], which rely on multiple datasets with varying camera parameters, we aim to learn this mapping using data acquired with a single camera system $\bm{\mathcal{C}}_{\text{Train}}$.
@@ 3.1 Depth Cues
For ground vehicles, a fixed position of the camera system relative to the ground introduces an additional perspective geometry constraint – known depth of the ground plane, which is constant for all images acquired by the same camera setup. For each pixel $(u,v)$, depth $d$ is calculated from the intersection of the camera rays $d\,\mathbf{K}^{-1}[u,v,1]^{T}$ and the ground plane $(^{\mathcal{C}}_{\mathcal{E}}\mathbf{R}\mathbf{n}_{\mathcal{E}})^{T}\mathbf{p}_{\mathcal{C}}+h=0$, with $\mathbf{n}_{\mathcal{E}}=[0,0,1]^{T}$ being the upward ground normal. This constraint enables depth regression via object’s ground contact point:
$d$ $=\frac{-h}{(^{\mathcal{C}}_{\mathcal{E}}\mathbf{R}\mathbf{n}_{\mathcal{E}})^{T}\mathbf{K}^{-1}[u,v,1]^{T}}$ (3)
$=\frac{f_{y}h}{\left(H-c_{y}-y\right)\cos(\theta)-f_{y}\sin(\theta)},$
where y=Hvy=H-v represents vertical image position of the ground projection, as visualized in Fig. 2. Compared to the object size cue, which is entirely semantic, this cue implicitly encodes the known perspective geometry constraints of the scene, thereby inducing a valuable prior to the ill-posed MDE problem. However, it is highly reliant on accurate and consistent camera calibration.
@@ 3.2 Canonical Transforms
Metric3D [49] proposed a canonical transform disentangling the predicted depth and camera parameters, enabling the use of object size cue for arbitrary camera systems. To achieve this, depth is predicted in canonical space $\mathbf{C_{F}}$ with a canonical focal length $f_{c}$, and post-processed to metric depth $\mathbf{D_{F}}$ by applying Focal Canonical Transform $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$, where $\mathbf{D_{F}}=\mathrm{FCT}_{\bm{\mathcal{C}}}(\mathbf{C_{F}})=\frac{f_{y}}{f_{c}}\mathbf{C_{F}}$, and $\bm{\mathcal{C}}=\{f_{y}\}$.
Besides the object size cue, in this paper we propose to leverage the vertical image position cue, effectively encoding the configuration of the scene’s perspective geometry. Unfortunately, this cue also induces entanglement between camera parameters and depth. To enhance cross-dataset generalization, we introduce an additional canonical transform, which enables consistent exploitation of the ground plane prior. Under this formulation, the model estimates a dense map $\mathbf{C_{Y}}$ representing vertical image position of the ground projection point. Metric depth $\mathbf{D_{Y}}$ is then obtained via Vertical Canonical Transform $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$, where $\mathrm{VCT}_{\bm{\mathcal{C}}}(\mathbf{C_{Y}})$ corresponds to mapping $y\mapsto d$ obtainable from Eq. 3. Therefore, this transformation requires knowledge of camera parameters $\bm{\mathcal{C}}=\{f_{y},c_{y},h,\theta\}$. Crucially, the proposed canonical space remains consistent under variations of camera parameters, facilitating generalization to diverse scenarios.
The proposed $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ is related to the depth of the ground plane, frequently used to enhance MDE accuracy in autonomous driving scenarios [46, 4, 41]. However, these methods exploit the ground-plane constraint exclusively for road pixels, resulting in limited generalization capability.
@@ 3.4 Geometric Augmentation
Contrary to zero-shot methods trained on multiple datasets [34, 16, 3, 49], we aim to achieve similar generalization capabilities by training on a single camera system. However, without geometric diversity induced by varying camera parameters, model can simply disregard our proposed canonical transformations. Therefore, we apply geometric augmentations, namely random cropping and resizing, to simulate different camera perspectives. Crucially, compared to [34, 3, 49], we train our model with multiple resolutions, enabling generalization for arbitrary image sizes.
@@ 3.5 Optimization
[…]
@@ 4 Experiments
Datasets. We use a diverse set of autonomous driving datasets, each collected with a different camera setup. Specifically, we utilize an aggregation of KITTI [13], DDAD [15], DrivingStereo [44], Waymo [39] and Argoverse Stereo [6] datasets. Due to the missing and sometimes inaccurate information, we calibrate extrinsic parameters, i.e., camera height and pitch, which are required for our canonical representation. Further details are provided in the supplementary materials.
@@
Evaluation Details. We use established depth metrics: absolute relative error ($\mathrm{A.Rel}$), root mean squared error ($\mathrm{RMS}$), root mean squared log error ($\mathrm{RMS}_{\log})$ and accuracy $\mathrm{\delta}_{i}$ under a threshold $1.25^{i}$. Depth is evaluated in the metric range of [0 m, 80 m].
@@ 4.1 Evaluation of Zero-shot Accuracy
Comparison with State-of-the-Art. We perform extensive evaluation, comparing GVDepth with SotA domain-dependent and domain-agnostic methods. Results for KITTI and DDAD datasets are shown in Tab. 1. Both supervised [33, 50] and self-supervised [14, 53] domain-dependent methods perform well on datasets matching their training domain but overfit to environmental and geometric biases, leading to reduced accuracy in KITTI $\rightarrow$ DDAD transfer. Relative depth estimation models exhibit greater adaptability across domains; however, they require ground-truth scale and shift alignment, limiting their practical applicability. Recent zero-shot metric methods, such as ZeroDepth, UniDepth, and Metric3D, demonstrate strong generalization capabilities but rely on extensive data acquisition and large-scale training, which is not always feasible. Furthermore, UniDepth and Metric3D suffer accuracy degradation when inferring on resolutions lower than their high-resolution training data. While resizing and padding can address this, these adjustments increase computational demands, limiting their suitability for resource-constrained systems. GVDepth delivers comparable results while being adaptable to arbitrary camera setups, aspect ratios, and image resolutions. Depending on the train-test dataset combination, our method ranks first or second across most metrics, despite training on only a single dataset and using a fraction of the data required by other zero-shot MDE methods. This is particularly impressive, as each dataset features unique environments and camera setups, demonstrating that our model relies on domain-agnostic perspective geometry cues rather than domain-specific semantic cues.
@@
Table 2: Comparison with GEDepth [46]. Vertical – depth regression via Vertical Canonical Transform - $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Fusion – depth regression with model as designed in Fig. 3. All models are trained with equivalent setup and model complexity on KITTI.
Testing GEDepth[46] Vertical Fusion
dataset $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$
KITTI
KITTI 5.68 95.4 5.70 95.5 5.67 95.7
DStereo 18.25 66.7 10.43 87.3 10.24 87.4
Waymo 22.11 69.1 13.42 78.6 14.34 78.4
Argo 19.27 52.8 14.72 64.7 10.45 84.8
DDAD 19.46 61.2 14.26 73.6 11.81 82.5
Our proposed $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ implicitly encodes depth of the ground plane, which has been utilized in [41, 46, 4] to improve MDE accuracy in autonomous driving scenes. However, as demonstrated in Tab. 2, $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ significantly outperforms GEDepth [46], with the performance gap further widened after fusion with $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. As we will show, the primary cause of such discrepency is inability of GEDepth to leverage vertical image positions and disentangle perspective geometry for pixels located outside the road surface. In the Supp. materials we provide a comprehensive analysis and critique of existing models that rely on ground plane depth.